# Application of Transformers for Nonlinear Channel Compensation in Optical Systems

### v2

@@ 3.3 Mask
The attention matrix can be masked for variety of reasons, for example to introduce some properties into the model such as preserving the auto-regressive property. Masking can also be used to reduce the complexity of attention calculations ($A$ in Eq.  (3)). One approach is to make the attention matrix sparse so that the zero elements are not computed or stored [23, 24]. Note that computing attention is a complexity bottleneck in the vanilla Transformer for long sequences. In this work, we propose a method to make the matrix $A$ sparse according to a physic-informed mask from the perturbation theory. As briefly discussed in Section 2, the optical field can be represented as the solution of a linear term plus a nonlinear perturbation term in symbol domain. The perturbed term can be represented as the weighted sum of symbol triplets assuming pulse spreading (memory) that is much greater than the symbol duration due to the accumulated dispersion impact [5]. The perturbation term for the target symbol at the middle of sequence can be simplified as:
@@
where $P_{0}$ is the optical launch power, $A_{.,x/y}$ is the sequence of complex transmitted symbols, and $C_{m,n}$ is the nonlinear perturbation coefficient.
Specifically, the perturbation coefficients $C_{m,n}$ show which combinations of symbols are more important to estimate the nonlinear interference of the target symbol. Therefore, we can use this two-dimensional relations inside our attention matrix as a mask to reduce the computational complexity. Considering the hyperbolic form of the perturbation coefficients, we can use a two-dimensional link-independent criteria as a mask by selecting a subset of elements in attention matrix. As an example, we use the following relation as in [6] for the numerical result:
@@
Since the Transformer also needs to determine the interaction between symbols, based on the above equation, we propose only using the elements of attention matrix $A_{m,n}$ that satisfy the $m$ and $n$’s relation in Eq. (8), where $m$ and $n$ are row and column indices, respectively, and make all the other elements zeros. To implement this approach in the individual symbol processing, in the simulation, we add a mask matrix, $M^{indv}$ to $QK^{T}/\sqrt{d_{k}}$. The matrix, $M^{indv}$, has the same size as $QK^{T}$ and is generated such that the elements that satisfy the relationship between $m$ and $n$ in Eq. (8) are filled by zeros and those that do not are filled with the negative infinity. This guaranties that the elements in attention matrix that do not meet Eq. (8) are set to zeros since the softmax operator maps the negative infinity to zero. The elements that their indices meet Eq. (8) are not affected by the mask since adding a zero to an element has no effects on the softmax’s output.
In summary, with the assumption that computation of the attention matrix can be implemented such that the computations are only carried out for the non-zero values in the attention matrix (corresponding to zero values in the proposed mask), we can reduce the computational/space complexity of attention to $O(\hat{N}d_{k})$ where $\hat{N}$ is the number of non-zero attention elements.
At first, we present the mask generation for an individual target symbol (block size of one). The algorithm for generating $M^{indv}$ is shown in Alg. (1). Depending on the values of tap size $t$ and mask hyperbolic parameter $\rho$, the number of non-zero elements in the mask varies and consequently the number of operations needed for computing the attention matrix changes. Figure 4 shows the attention mask for the individual symbol processing for different values of $t$ and $\rho$.
Algorithm 1 Algorithm for mask generation of an individual target symbol.
@@
Figure 4: Attention masks for the cases where the output error only computed for the individual target symbol. The red area shows the values that are zeros (unmasked) and the blue area shows the values which are negative infinity (masked). (a) $t=32$ and $\rho=2.6$, (b) $t=64$ and $\rho=2.6$, (c) $t=64$ and $\rho=1.3$, and (d) $t=64$ and $\rho=0.4$. The ratio of number of unmasked (zero) elements to the total number of elements is 0.31, 0.19, 0.10 and 0.04 for figure (a), (b), (c), and (d), respectively.
Next, we discuss the mask approach for the processing of a block of target symbols. The method we propose here is based on computing a union-like combination of the individual symbol’s mask. The algorithm for computing the mask for the block-processing approach is shown in Alg. (2). Note that the size of the block mask matrix, $M^{block}$, is $(\ell+b)\times(\ell+b)$ and the size for individual symbol’s mask, $M^{indv}$, is $(\ell+1)\times(\ell+1)$, where individual mask is computed using Alg. (1). Two block masks are depicted in Fig. 5 where the block mask for a model with $t=64,b=128,\rho=2.6$ is shown in Fig. 4(a) and Fig. 4(b) shows a block mask for a model with $t=64,b=128,\rho=0.4$. As can be seen from this figure, the choice of $\rho$ does not significantly alter the mask which is not surprising since the block mask is generated by stacking individual masks on each others. The ratios of unmasked (zero) elements to the total elements are $0.34$ and $0.31$ for $\rho=2.6$ and $\rho=0.4$, respectively.
Algorithm 2 Algorithm for mask generation of a block of target symbols.
@@
$M^{indv}$ = generate mask for one symbol using Alg. 1
@@
Figure 5: Masks for blocks: (a) shows a block mask with $t=64,b=128,\rho=2.6$ while (b) depicts a block mask with $t=64,b=128,\rho=0.4$. (c) shows the mask with $t=64,b=4096,\rho=2.6$. The blue area shows the elements with negative infinity values and the red shows zero elements. (d) shows the ratio of number of unmasked (zero) elements to the total number of elements versus block size.
For large block sizes, this ratio substantially decreases. Fig. 4(c) shows the impact of increasing the block size to 4096. In this case, only $3\%$ of elements of the attention matrix are preserved which results in a large complexity reduction. Finally, Fig. 4(d) shows the ratio of unmasked (zero) elements to the total number of elements as a function of block size. As it can be seen, the ratio first increases and then decreases as the block size increases. It should be noted that the choice of block-size needs to be optimized according to the system requirements in terms of throughput and latency, and hence, it cannot be increased too much due to increasing required memory and computational complexity of Transformers. A detailed discussion on the impact of block size is given in the Appendix. Furthermore, to give some numerical examples of the effect of mask on the softmax input, several two-dimensional heat maps of $(QK^{T})/\sqrt{d_{k}}$ for Transformer with and without masks are shown in Appendix A.
3.4 Using Neighbors at the Output Layer of Transformer
For computing nonlinear distortion from the output of the last layer of Transformer, we employ an MLP as explained in Section 3.1. To compute each nonlinear distortion $s_{i}$, one option is to give the corresponding Transformer-generated representation, $r_{i}$, to the MLP. However, due to the dependency of nonlinear estimates, we observe that if for each symbol $s_{i}$, we also provide few representations around $r_{i}$ to the MLP, so it can compute $s_{i}$ more accurately. We call this approach window processing and while it can increase the model’s performance, it does not increase the computational complexity substantially for small and moderate window sizes.
@@ 3.4 Using Neighbors at the Output Layer of Transformer
4 Numerical Results
4.1 System Model
@@
Here, we focus on two cases with single-channel coherent system. The first one is operating at 32 Gbaud 16QAM modulation where the link consists of 40 spans of standard single-mode fiber (SSMF). The second case uses 42 Gbaud 64QAM modulation format with 12 spans of SSMF. Each span has a length of 80km followed by an optical amplifier with $6$ dB noise figure. The signals are pulse shaped by a root-raised cosine function with a roll-off factor of $1/16$. Furthermore, we consider a symmetric dispersion map, in which half of total dispersion is digitally pre-compensated at the transmitter side. Training is performed at $2$ dBm and $3$ dBm launch powers for the two cases, respectively. This is close to the optimal launch power for each case when DBP at two samples-per-symbol and 1 to 3 steps-per-span are employed to benchmark these results.
@@
Figure 6: System model with nonlinear equalizer at receiver side.
4.2 Numerical Setup
@@
Standard Rx-DSP output for $2^{19}$ and $2^{18}$ symbols were used in the training and evaluation of ANN models, respectively. A permuted congruential generator (PCG64) with different seeds for training and evaluation stages was employed. Models are trained for a mini-batch size of 512 using Adam optimizer with a learning rate scheduler that includes a warm-up stage, and a mean-squared-error loss function. Early stopping is used to stop the training if the model performance has not improved over 100 epochs. Note that the proposed ANN-NLC equalizers estimate the nonlinear distortions in one polarization. However, due to the symmetric nature of signal propagation in fiber medium, it has been shown that the same model can be used to generate the nonlinear error estimates for the other polarization by simply swapping input signals. This enables efficient learning of a generalized model that performs on both polarizations.
@@
Parameters and components in a Transformer architecture can be varied in order to explore performance and computational complexity of different realizations. We perform a wide grid search over hyper-parameters of the Transformer and various components consisting of hundreds of models. Among these parameters, we studied the impact of following hyper-parameters: tap size, $d_{model}$ which we will call hidden size, key size, number of heads, number of encoder layers, number of outputs for the first layer of point-wise FFN which is called FFN’s hidden size, presence or absence of mask, and output layer window size. We implemented the search in two steps: At first, for each hyper-parameter we ran a specific grid search where we selected a wide range for that specific hyper-parameter while only few values for the remaining parameters were included. In the second step, we chose some of the best values from the first run and performed a grid search on the combination of those values. As discussed earlier, the type of network that generates the embeddings is important. Hence, we carried out a separate grid-search on the impact of block size and embedding generator networks on a smaller grid which is presented in the Appendix. Table 1 shows the summary for selected values of the grid search in the second step. We employed separate search spaces for 16QAM and 64QAM cases tailored to their propagation mediums.
Table 1: Transformer’s hyper-parameters and their values used for the grid search.
  Hyper-parameter Grid Search Values for 16QAM Grid Search Values for 64QAM
@@
4.3 Analysis of the Performance versus Complexity Trade-off for 16QAM Setup
@@
The numerical results for hyper-parameters grid search in the 16QAM case is given in Fig. 7. This figure shows the scatter plot for the performance, represented by the quality metric Q, versus complexity, represented by RMPS, for all model variations obtained from sweeping over hyper-parameters in Table 1. We use blue and red colored dots to show the results for Transformers with and without mask, respectively. Additionally, the blue and red curves illustrate the envelopes associated to the best performing Transformers with and without mask, respectively. All Transformer-NLCs are designed for a block size of 128 symbols. As seen from the figure, Transformers can achieve impressive nonlinear compensation performance as the best models in 16QAM setup have improved Q from $6.7$ dB (linear case) to more than $8.8$ dB, which is $2.1$ dB improvement at the training launch power of $2$ dBm. In addition, even at the lower complexities where the performance has been decreased, they still retain a good nonlinear compensation capability. Moreover, using a mask particularly boost the performance at lower complexity regions compared to the Transformers without mask while there is no degrading impact on the performance at higher complexities.
@@
Figure 7: Performance vs. complexity trade-off for the 16QAM setup for the grid search shown in Table 1. The models were trained and evaluated at $2$ dBm launch power. The red dots show the results for the models that did not use the proposed mask while blue triangles represent the ones with the proposed mask.
To better illustrate performance results and have a clearer picture for the impact of the proposed masks, an ensemble of models with and without masks are selected for further analysis. These models are chosen from different regions on the envelope of performance versus complexity trade-off graph. Table 2 contains the hyper-parameters for the selected models in 16QAM setup. The performance versus launch power for the three selected models at different computational complexity regions is depicted in Fig. 8. Note that the models are trained at 2 dBm launch power and evaluated at different powers by scaling the estimated nonlinear distortion according to Eq. (2). We observe that the selected models demonstrate good generalization and can provide nonlinear performance gain over a wide range of launch powers. Additionally, compared to linear compensation, the best performing model improves the optimal launch power by 1.5 dB while the maximum Q-value at the optimal launch power is also boosted by 1.5 dB (from 7.3 dB at 0.5 dBm to 8.8 dB at 2 dBm). Table 3 illustrates the summary of results for the above-selected models at 2 dBm launch power as an example. The improvement of performance-complexity trade-off as a result of employing the attention mask is clear for the model in region 3, corresponding to the least complexity region.
@@
It is worthwhile to note that at mid and high complexity regions, the proposed mask helps the training process by pointing to the most important attention positions and hence, it saves in computational complexity for a given well trained model (as Region 1 and 2 in Table 3 and Fig. 8). On the other hand, since generally there is enough model capacity to learn parameters in those regions, the unmasked network can also learn those important relationships. Therefore, using mask reduces the complexity with minimal changes in performance. However, that is not the case for the lower complexity region. In fact, directing limited resources toward important model parameters provides a learning advantage here where one can see the performance boost in this region. In other words, with under-parameterized models in low complexity region, the mask guidance generally results in a more efficient training process. Specifically, for the fixed model in Region 3, one can see a better performance while the complexity is still reduced by using the proposed mask.
Moreover, performance results for DBP at two samples-per-symbol with different steps-per-span (StpS) are presented to benchmark the proposed Transformers. Note that, the complexity comparison of DBP is not performed here since it involves implementation obstacles beyond multiplication complexity. In fact, DBP faces serious challenges for implementation in coherent transceivers including higher required over-sampling rate, the need for more number of steps-per-spans with increasing symbol rate to maintain the performance, and also use of multiple (inverse) fast Fourier transform (FFT) modules which impacts the linear equalization data path. Hence, DBP serves as performance upper-bound at two samples-per-symbol where with no other component or channel impairments involved here, it can efficiently reverse the simulated nonlinear channel.
Table 2: Hyper-parameters for the selected models at three different performance vs. complexity regions for 16QAM setup.
  Hyper-parameter Region 1 Region 2 Region 3
tap size 96 64 16
hidden size 96 64 16
key size 64 32 16
number of heads 4 4 4
number of encoder layer 3 2 2
FFN’s hidden size 64 32 32
window size 15 7 7
Figure 8: Performance of the three selected 16QAM models of Table 2 at various launch powers. The results of the linear compensation and DBP with 1 and 2 steps-per-span are also shown.
Table 3: Summary of evaluated results at 2 dBm launch power for selected models with 16QAM.
@@
Q $\mathbf{Q_{NN}}$ RMPS (K) $\mathbf{Q_{NN}}$ RMPS (K)
Region 1 6.67 8.82 608 8.76 404
Region 2 6.67 8.54 129 8.51 81
Region 3 6.67 7.31 22 7.68 9
4.4 Analysis of the Performance versus Complexity Trade-off for 64QAM Setup
@@
As for the previous case, we also performed a grid search for 64QAM setup. All models were trained with a block size of 128 symbols. Fig. 9 shows the scatter plot for the performance versus complexity, for all model variations by sweeping over the hyper-parameters for 64QAM in Table 1. In the figure, blue and red colored dots are used to show the results for Transformers with and without mask, respectively, and blue and red lines are employed to depict the envelopes associated to the best performing Transformers with and without mask, respectively.
We observe that the Transformers provide an excellent nonlinear compensation performance as the best models have improved Q from 6.5 dB (linear case) to 8.17 dB, a 1.67 dB improvement at the training launch power of 3 dBm. Also in this setup, the attention mask boosts the performance at lower complexity regions with almost no loss in general.
@@
Figure 9: Performance vs. complexity trade-off graph of the 64QAM setup for the grid search shown in Table 1. Models were trained and evaluated at 3 dBm launch power. Red dots (blue triangles) represent models with (without) the proposed mask, respectively.
Similar to the previous case, we selected three models as examples from different regions close to the envelope of the performance versus complexity trade-off curve. Table 4 shows the hyper-parameters for the selected model in each region. Fig. 10 shows the launch-power-sweep performance for the three selected models as well as DBP results at two samples-per-symbol. Note that Transformer-NLC models are trained at 3 dBm launch power and evaluated at different powers by adjusting the scale according to Eq. (2). This figure shows that the selected models achieve good generalization and provide great nonlinear performance gains over a wide range of launch powers. As it can be seen, compared to the linear compensation, the best performing model improves the optimal launch power by 1.5 dB while the maximum Q-value at the optimal launch power is also increased by 1.26 dB (from 6.91 dB at 1.5 dBm to 8.17 dB at 3 dBm). Finally, Table 5 shows the summary of results for the above-selected regions at 3 dBm launch power as an example. As it can be seen, in Region 1 and 2, the model with the mask performs as well as the model without a mask while has a lower RMPS. In Region 3, the use of attention mask reduced the complexity while improved the performance at the lowest complexity region.
Table 4: Hyper-parameters for the selected models at three different performance vs. complexity regions for 64QAM setup.
  Hyper-parameter Region 1 Region 2 Region 3
tap size 64 32 8
hidden size 96 64 16
key size 32 32 16
number of heads 4 4 4
number of encoder layer 3 1 1
FFN’s hidden size 64 64 64
window size 11 11 7
Table 5: Summary of evaluated results at 3 dBm launch power for selected 64QAM models.
  No Mask Mask with $\bm{\rho=2.6}$
Q $\mathbf{Q_{NN}}$ RMPS (K) $\mathbf{Q_{NN}}$ RMPS (K)
Region 1 6.50 8.16 268 8.17 196
Region 2 6.50 7.9 50 7.84 35
Region 3 6.50 7.02 11 7.20 5
Figure 10: Performance of selected 64QAM models for the three different regions of Table 5 at various launch powers . The results of the linear compensation and DBP with 1 and 3 steps-per-span are also shown.
4.5 A comparison to an LSTM-based equalizer

### v3

@@ III-C Mask
The attention matrix can be masked for variety of reasons, for example to introduce some properties into the model such as preserving the auto-regressive property. Masking can also be used to reduce the complexity of attention calculations ($A$ in Eq.  (3)). One approach is to make the attention matrix sparse so that the zero elements are not computed or stored [26, 27] since computing attention is a complexity bottleneck in the vanilla Transformer for long sequences. In this work, we propose a method to make the matrix $A$ sparse according to a physic-informed mask from the perturbation theory. As briefly discussed in Section II, the optical field can be represented as the solution of a linear term plus a nonlinear perturbation term in symbol domain. The perturbed term can be represented as the weighted sum of symbol triplets assuming pulse spreading (memory) that is much greater than the symbol duration due to the accumulated chromatic dispersion in the fiber [5]. The perturbation term for the target symbol at the middle of sequence can be represented as:
@@
where $P_{0}$ is the optical launch power, $A_{.,x/y}$ is the sequence of complex transmitted symbols, and $C_{m,n}$ is the nonlinear perturbation coefficient. Specifically, the perturbation coefficients $C_{m,n}$ show which combinations of symbols are more important to estimate the nonlinear interference of the target symbol. Therefore, we can use this two-dimensional relations inside our attention matrix as a mask to reduce the computational complexity.
Considering the hyperbolic form of the perturbation coefficients, we can use a two-dimensional link-independent criteria as a mask by selecting a subset of elements in attention matrix. As an instance, we use the following relation for the numerical result as in [6]:
@@
Since the Transformer also needs to determine the interaction between symbols, based on the above equation, we propose only to use the elements of attention matrix $A_{m,n}$ that satisfy the $m$ and $n$’s relation in Eq. (8), where $m$ and $n$ are row and column indices, respectively, and make all the other elements zeros. To implement this approach in the individual symbol processing, in the simulation, we add a mask matrix, $M^{indv}$ to $QK^{T}/\sqrt{d_{k}}$. The matrix, $M^{indv}$, has the same size as $QK^{T}$ and is generated such that the elements that satisfy the relationship between $m$ and $n$ in Eq. (8) are filled by zeros and those that do not are filled with the negative infinity. This guaranties that the elements in attention matrix which do not satisfy Eq. (8) are set to zeros since the softmax operator maps the negative infinity to zero. The elements that their indices meet Eq. (8) are not affected by the mask since adding a zero to an element has no effects on the softmax’s output. In summary, with the assumption that computation of the attention matrix can be implemented such that the computations are only carried out for the non-zero values in the attention matrix (corresponding to zero values in the proposed mask), we can reduce the computational/space complexity of attention to $O(\hat{N}d_{k})$ where $\hat{N}$ is the number of non-zero attention elements.
At first, consider the mask generation for an individual target symbol (block size of one). The algorithm for generating $M^{indv}$ is shown in Alg. (1). Depending on the values of tap size $t$ and mask hyperbolic parameter $\rho$, the number of non-zero elements in the mask varies and consequently the number of operations needed for computing the attention matrix changes. Figure 4 shows the attention mask for the individual symbol processing for different values of $t$ and $\rho$.
Algorithm 1 Mask generation for an individual target symbol.
@@
Figure 4: Attention masks for the cases where output error only computed for the individual target symbol. The red area shows the values that are zeros (unmasked) and the blue area shows the values which are negative infinity (masked). (a) $t=32$ and $\rho=2.6$, (b) $t=64$ and $\rho=2.6$, (c) $t=64$ and $\rho=1.3$, and (d) $t=64$ and $\rho=0.4$. The ratio of number of unmasked (zero) elements to the total number of elements is 0.31, 0.19, 0.10 and 0.04 for figure (a), (b), (c), and (d), respectively.
Next, we discuss the mask approach for the processing of a block of target symbols. The method we propose here is based on computing a union-like combination of the individual symbol’s mask. The algorithm for computing the mask for the block-processing approach is shown in Alg. (2). Note that the size of the block mask matrix is $M^{block}$, is $(\ell+b)\times(\ell+b)$ while the size for individual symbol’s mask was $M^{indv}$, is $(\ell+1)\times(\ell+1)$. For two different values of $\rho$, block masks are depicted in Fig. 5 for a model with $t=64,b=128,\rho=2.6$ in Fig. 5(a) and another model with $t=64,b=128,\rho=0.4$ in Fig. 5(b), respectively. As can be seen from this figure, the choice of $\rho$ does not significantly alter the mask which is not surprising since the block mask is generated by stacking individual masks on each others. The ratios of unmasked (zero) elements to the total elements are $0.34$ and $0.31$ for $\rho=2.6$ and $\rho=0.4$, respectively.
Algorithm 2 Mask generation for a block of target symbols.
@@
$M^{indv}$ = generate mask for one symbol using Algorithm 1
@@
Figure 5: Masks for blocks: (a) shows a block mask with $t=64,b=128,\rho=2.6$ while (b) depicts a block mask with $t=64,b=128,\rho=0.4$. (c) shows the mask with $t=64,b=4096,\rho=2.6$. The blue area shows the elements with negative infinity values and the red shows zero elements. (d) shows the ratio of number of unmasked (zero) elements to the total number of elements versus block size for several values of $\rho$s.
For large block sizes, this ratio substantially decreases. Fig. 5(c) shows the impact of increasing the block size to 4096. In this case, only $3\%$ of elements of the attention matrix are preserved which results in a large complexity reduction. Finally, Fig. 5(d) shows the ratio of unmasked (zero) elements to the total number of elements as a function of block size for several values of $\rho$s. As the figure shows when the block size is increased the value of $\rho$ has less effects on the ratio of the number of unmasked elements to the total number of elements In fact, the ratio first increases and then decreases as the block size increases. It should be noted that the choice of block-size needs to be optimized according to the system requirements in terms of throughput and latency, and hence, it cannot be increased too much due to increasing required memory and complexity of Transformers. A detailed discussion on the impact of block size is given in [28]. To give some numerical examples of the effect of mask on the softmax input, several two-dimensional heat maps of $(QK^{T})/\sqrt{d_{k}}$ for Transformer with and without masks are shown in Appendix I.
III-D Using Neighbors at the Output Layer of Transformer
For computing nonlinear distortion from the output of the last layer of the Transformer, we employ an MLP as discussed in Section III-A. To compute each nonlinear distortion $s_{i}$, one option is to give the corresponding Transformer-generated representation, $r_{i}$, to the MLP. However, due to the dependency of nonlinear estimates, we observe that if for each symbol $s_{i}$, we also provide few extra representations around $r_{i}$ (the neighbors) to the MLP, it can compute $s_{i}$ more accurately. We call this approach window processing and while it can increase the model’s performance, it does not increase the computational complexity substantially for small and moderate window sizes. Specifically, the selected window of representations are fed into an MLP module consists of a linear layer ($\in R^{W\cdot d_{model}\times 2}$), a leaky ReLU nonlinear activation function with negative slope of $0.2$, a second linear layer ($\in R^{2\times 10}$), a leaky ReLU nonlinear activation function with negative slope of $0.2$, and a third linear layer ($\in R^{10\times 2}$). The last layer has two outputs that generates the estimated nonlinear distortions corresponding to the target symbols.
@@ III-D Using Neighbors at the Output Layer of Transformer
IV Numerical Results
IV-A System Model
@@
To study the performance versus complexity trade-off, we focus on two cases with single-channel coherent system. The first one is operating at 32 Gbaud 16QAM modulation where the link consists of 40 spans of standard single-mode fiber (SSMF). The second case uses 42 Gbaud 64QAM modulation format with 12 spans of SSMF. Each span has a length of 80km followed by an optical amplifier with $6$ dB noise figure. The signals are pulse shaped by a root-raised cosine function with a roll-off factor of $1/16$. Furthermore, we consider a symmetric dispersion map, in which half of total dispersion is digitally pre-compensated at the transmitter side. Training is performed at $2$ dBm and $3$ dBm launch powers for the two cases, respectively. This is close to the optimal launch power for each case.
@@
Figure 6: System model with nonlinear equalizer at the receiver side.
IV-B Numerical Setup
@@
Standard Rx-DSP output for $2^{19}$ and $2^{18}$ symbols were used in the training and evaluation of ANN models, respectively. A permuted congruential generator (PCG64) with different seeds for training and evaluation stages was employed. Models are trained for a mini-batch size of 512 using Adam optimizer with a learning rate scheduler that includes a warm-up stage, and a mean-squared-error loss function. Early stopping is used to stop the training if the model performance has not improved for over 100 epochs. Note that the proposed ANN-NLC equalizers estimate the nonlinear distortions in one polarization. However, due to the symmetric nature of signal propagation in fiber medium, it has been shown that the same model can be used to generate the nonlinear error estimates for the other polarization by simply swapping input signals. This enables efficient learning of a generalized model that performs on both polarizations. Furthermore, the output layer can be slightly modified in order to generate the nonlinear estimates for both polarizations at the same time.
@@
Parameters and components in a Transformer architecture can be varied in order to explore performance and computational complexity of different realizations. We perform a wide grid search over hyper-parameters of the Transformer and various components consisting of hundreds of models. Among these parameters, we studied the impact of following hyper-parameters: tap size, $d_{model}$ which we will call hidden size, key size, number of heads, number of encoder layers, number of outputs for the first layer of point-wise FFN which is called FFN’s hidden size, presence or absence of mask, and output layer window size. We implemented the search in two steps: At first, for each hyper-parameter we ran a specific grid search where we selected a wide range for that specific hyper-parameter while only few values for the remaining parameters were included. In the second step, we chose some of the best values from the first run and performed a grid search on the combination of those values. As discussed earlier, the type of network that generates the embeddings is important. Hence, we carried out a separate grid-search on the impact of block size and embedding generator networks on a smaller grid where the details are presented in [28]. Table I shows the summary of selected values for the grid search in the second step. We have employed separate search spaces for 16QAM and 64QAM cases tailored to their propagation mediums.
TABLE I: Transformer’s hyper-parameters used for the grid search.
  Hyper-parameter Grid-Search Values (16QAM) Grid-Search Values (64QAM)
@@
IV-C Performance vs. Complexity Trade-off for 16QAM Setup
@@
The numerical results for hyper-parameters grid search in the 16QAM case is given in Fig. 7. This figure shows the scatter plot for the performance, represented by the quality metric Q, versus complexity, represented by RMPS, for all model variations obtained from sweeping over hyper-parameters in Table I. We use blue and red colored dots to show the results for Transformers with and without mask, respectively. Additionally, the blue and red curves illustrate the envelopes associated to the best performing Transformers with and without mask, respectively. All Transformer-NLCs are designed for a block size of 128 symbols. As seen from the figure, Transformers can achieve impressive nonlinear compensation performance as the best models in 16QAM setup have improved Q from $6.7$ dB (linear case) to more than $8.8$ dB, which is $2.1$ dB improvement at the training launch power of $2$ dBm. In addition, even at the lower complexities where the performance has been decreased, they still retain a good nonlinear compensation capability. Moreover, using a mask particularly boost the performance at lower complexity regions compared to the case without mask while there is no degrading impact on the performance at higher complexities.
@@
Figure 7: Performance vs. complexity trade-off for the 16QAM setup with grid search shown according to Table I. The models were trained and evaluated at $2$ dBm launch power. Red dots show the results for models that did not use any mask while blue triangles represent the ones with the proposed mask.
@@
It is worthwhile to note that at mid and high complexity regions, the proposed mask helps the training process by pointing to the most important attention positions and hence, it can save in computational complexity for a given well trained model. However, for the lower complexity region, there is a performance advantage as well. In fact, directing limited resources toward important model parameters provides a learning advantage here where one can see the performance boost in this region. In other words, with under-parameterized models in low complexity region, the mask guidance generally results in a more efficient training process. Table II shows the performance and complexity of three models as examples of different performance versus complexity regions from Fig. 7.
TABLE II: Impact of applying mask on performance and complexity of three selected models at 2 dBm launch power for 16QAM.
@@
Model Q(dB) QNN(dB) RMPS(K) QNN(dB) RMPS(K)
A 6.67 8.82 608 8.76 404
B 6.67 8.54 129 8.51 81
C 6.67 7.31 22 7.68 9
Finally, note that a deeper study on impact of different hyper-parameters on performance versus complexity of the proposed Transformer-NLC has been presented in [28]. The details are omitted here for brevity.
IV-D Performance vs. Complexity Trade-off for 64QAM Setup
@@
As for the previous case, we also perform a grid search for 64QAM setup. All models are trained with a block size of 128 symbols. Fig. 8 shows the scatter plot for the performance versus complexity, for all model variations by sweeping over the hyper-parameters for 64QAM in Table I. In the figure, blue and red colored dots are used to show the results for Transformers with and without mask, while the blue and red lines are employed to depict the envelopes associated to the best performing Transformers with and without mask, respectively. We observe that the Transformers provide an excellent nonlinear compensation performance as the best models have improved Q from 6.5 dB (linear case) to 8.17 dB, a 1.67 dB improvement at the training launch power of 3 dBm. Also in this setup, the attention mask boosts the performance at lower complexity regions with almost no loss in general.
@@
Figure 8: Performance vs. complexity trade-off graph of the 64QAM setup for the grid search shown in Table I. Models were trained and evaluated at 3 dBm launch power. Red dots (blue triangles) represent models with (without) the proposed mask, respectively.
IV-E Comparison to an LSTM-based equalizer